A speedboat moving at 30.0 m/s approaches a ship 100 m ahead. The pilot slows the boat with a constant acceleration of -3.50 m/s2. How long does it take the boat to reach the ship?
1
Expert's answer
2020-10-12T11:31:05-0400

Using "3^{rd}"  kinematics equation-

"s=ut+0.5at^{2}"

here,

s is the displacement

u is the initial velocity

a is the acceleration

t is time

Given : s = 100m , a = -3.50m"s^{-2}" , u = 30m"s^{-1}"

Putting the values-

"100=30t-1.75t^2"

"175t^2-3000t+10000=0"

"(t-4.531)(t-12.612)=0"

Here t = 12.612s is not possible. Hence, time taken by boat to reach the ship is 4.531s